# Audio-Agent: Leveraging LLMs For Audio Generation, Editing and Composition

### v1

@@ 3.5 Conditional Audio Generation
$\begin{split}\mathbf{Z}^{new}=\text{Softmax}(\frac{\mathbf{Q}\mathbf{K}_{txt}^{\top}}{\sqrt{d}})\mathbf{V}_{txt}+\text{Softmax}(\frac{\mathbf{Q}(\mathbf{K}_{vis})^{\top}}{\sqrt{d}})\mathbf{V}_{vis}\\ \text{where}\ \mathbf{Q}=\mathbf{Z}\mathbf{W}_{txt}^{q},\mathbf{K}_{txt}=\bm{c}_{txt}\mathbf{W}_{txt}^{k},\mathbf{V}_{txt}=\bm{c}_{txt}\mathbf{W}_{txt}^{v},\\ \mathbf{K}_{vis}=\bm{c}_{vis}\mathbf{W}_{vis}^{k},\mathbf{V}_{vis}=\bm{c}_{vis}\mathbf{W}_{vis}^{v}\end{split}$ (2)
@@
$L_{\text{simple}}=\mathbb{E}_{\bm{x}_{0},\bm{\epsilon},\bm{c}_{txt},\bm{c}_{vis},t}\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\big{(}\bm{x}_{t},\bm{c}_{txt},\bm{c}_{vis},t\big{)}\|^{2}.$ (4)
Compared to IP-Adapter Ye et al. (2023), our method introduces the video modality into audio generation. Furthermore, since the semantic tokens already incorporate temporal information of the video, we do not need to train an extra timestamp detection module as done by FoleyCrafter Zhang et al. (2024b) to achieve temporal alignment.
@@ 3.6 Implementation Details
For fine-tuning Gemma-2B-it, we set LoRA rank and alpha to be 64 with dropout to be 0.05. We separately train and fine-tune Gemma-2B-it, the projection layers and the cross-attention layers on the AVSync15 Zhang et al. (2024a) datasets. The training and evaluation are conducted on NVIDIA GeForce RTX 4090. Following Ye et al. (2023), we set the $\lambda$ to be 0.5 as default.
@@ 4.1 Training Datasets
For the TTA task, we evaluate our complex generation ability on AudioCaps Kim et al. (2019) dataset. We randomly choose either one caption from the test set or concatenate two of them together with the clause “followed by”. To better compare with other models, we limit our generation length to the standard 10 seconds. Following Xue et al. (2024), we randomly selected 20 captions from each category for the generation. Additionally, to demonstrate Audio-Agent’s ability to make inferences based on complex text conditions, we ask GPT to generate additional long event descriptions containing a series of events without explicitly describing the sound, such as “A man enters his car and drives away”. The number of complex captions is also 20. The baseline methods include AudioGen-v2-medium Kreuk et al. (2022), AudioLDM2-large Liu et al. (2024) and Auffusion Xue et al. (2024).
@@
We use AVSync15 for VTA task. AVSync15 is a curated dataset from VGGSound Sync Chen et al. (2021) that has 1500 high video-audio alignment pairs, which is ideal for training and demonstrating temporal alignment between video and audio. Same experiment setting as Zhang et al. (2024b) is used. To better facilitate evaluation, we include some audio generation results in the supplementary material.
4.2 Evaluation Metrics
The evaluation metrics are summarized as follows: For the VTA task, we use the Frechet audio distance (FAD) to evaluate audio fidelity. Additionally, we utilize the MKL metric Iashin & Rahtu (2021) and CLIP similarity Wu et al. (2022) for audio-video relevance. Furthermore, to evaluate the synchronization of the generated audio in the video-to-audio setting, we use the same evaluation metrics as CondFoleyGen Du et al. (2023), namely # Onset Accuracy, and Onset AP. For the TTA task, we use CLAP similarity Wu et al. (2023)
4.3 Evaluation and comparison
Audio-Agent outperforms other baseline methods on all TTA experiment settings, see Table 2. Additionally, our method outperforms the original Auffusion model by a significantly increasing margin as the text condition becomes longer and more complex. Specifically, we notice that with a longer text condition, AudioGen Kreuk et al. (2022), AudioLDM2 Liu et al. (2024) and Auffusion Xue et al. (2024) all exhibit missing out events. For example, if the text condition is multi-event such as “Pigeons cooing and bird wings flapping as footsteps shuffle on paper followed by motor sounds with male speaking”, all the baseline methods fail to generate the motor sound at the end of the audio clip during evaluation. However, our method avoids this problem by utilizing GPT-4 as a brain/coordinator for caption analysis and generation planning, offering more fine-grained distinctions between events.
We also notice a significant drop for all methods on complex captions, since none of these methods has been trained on this type of text condition. Still, we find this type of text condition more practical in the real world, since it does not require explicit descriptions of the characteristics of sound, but rather describes the scenario for sound generation, offering more flexibility for the sound producer. We attach some examples of complex results that we used for evaluation in Appendix A.3.
For the VTA task, our method achieves better visual-audio synchronization compared to other baseline methods, while subpar the current state-of-the-art method in terms of generation audio quality, presented in Tables 3 and 4. We consider this reasonable as most of the other baseline methods have been trained on multiple larger datasets.
Specifically, we find that the temporal connector may negatively affect the generated audio quality on a small scale. However, for the evaluation of synchronization, we noticed a significant improvement after the temporal connector was applied, especially for the Onset AP. Without explicit training of a timestamp detector, our method achieves a better performance in terms of onset Acc and Onset AP, see Figure 6 for illustration.
Figure 5: Comparison with baseline for TTA task. To demonstrate audio generation based on long complex text conditions, we ask the model to generate audio clips for 20 seconds. The text condition is drawn from the Two Captions category of Table 2: (A) A river stream of water flowing followed by typing on a computer keyboard; (B) A woman delivering a speech followed by a male speech and statics; (C) A vehicle engine revving then accelerating at a high rate as a metal surface is whipped followed by tires skidding followed by a door shutting and a female speaking; (D). Continuous white noise followed by a vehicle driving as a man and woman are talking and laughing; We can see that our method successfully generates multi-event audio at different times based on descriptions, while Auffusion mixes the generated audio.
@@ 4.3 Evaluation and comparison
AudioGen Kreuk et al. (2022) 49.34% 44.76% 23.98%
AudioLDM2 Liu et al. (2024) 47.04% 36.03% 23.33%
Auffusion Xue et al. (2024) 50.91% 45.90% 14.40%
@@
Figure 6: Comparison with baseline for VTA generation task. Compared to the baseline, the event occurrence is more explicit. Our method can produce audio that is more aligned and better synchronized with the input video.
Table 3: Quantitative evaluation on semantic alignment and audio quality. Specifically, Audio-Agent achieves on par performance versus state-of-the-art models in terms of Mean KL Divergence (MKL) Iashin & Rahtu (2021), CLIP Wu et al. (2022) and FID Heusel et al. (2017) on AVSync15  Zhang et al. (2024a).
@@
SpecVQGAN (Inception) Iashin & Rahtu (2021) 5.339 6.610 114.44
SpecVQGAN (ResNet) Iashin & Rahtu (2021) 3.603 6.474 75.56
Diff-Foley Luo et al. (2024) 1.963 10.38 65.77
Seeing and Hearing Xing et al. (2024) 2.547 2.033 65.82
FoleyCrafter Zhang et al. (2024b) 1.497 11.94 36.80
@@
Table 4: Quantitative evaluation in terms of temporal synchronization. We report onset detection accuracy (Onset ACC) and average precision (Onset AP) for the generated audios on AVSync Zhang et al. (2024a), which provides onset timestamp labels for assessment, following previous studies  Luo et al. (2024); Xie et al. (2024).
@@
SpecVQGAN(Inception) Iashin & Rahtu (2021) 16.81 64.64
SpecVQGAN(ResNet) Iashin & Rahtu (2021) 26.74 63.18
Diff-Foley Luo et al. (2024) 21.18 66.55
Seeing and Hearing Xing et al. (2024) 20.95 60.33
FoleyCrafter Zhang et al. (2024b) 28.48 68.14
@@ 4.4 Ablation Studies
We include our ablation study on different LoRA rank values during LLM fine-tuning, see Tables 5 and 6. We found that an increase in trainable parameters sometimes does not necessarily improve the result. Notwithstanding, for a fair comparison, we use the rank value of 60 across all metrics. Additionally during training, we found that the training of the cross-attention layer can converge within 20,000 steps. We notice that the loss curve is not a reliable indicator of the model’s performance. The model can achieve a good performance even when the loss curve remains flat.
Table 5: Ablation study on AVSync15 dataset with different LoRA rank for semantic alignment and audio quality. During experiments, we keep the value of alpha the same as the rank.
@@
Table 6: Ablation study on AVSync15 dataset with different LoRA rank in terms of temporal synchronization. During experiments, we keep the value of alpha the same as the rank.
@@ 5 Conclusion and Discussion
5.1 Limitation and Future Work
Our framework experiences a drop in performance when given complex text conditions for the TTA task, which is more severe in other baseline methods. We believe it is a worthwhile direction in the future for understanding long complex captions with improved fine-grained distinctions between multiple events. We may also utilize the LLM’s versatility involving audio captioning tasks and video captioning tasks. The above are worthwhile future directions to explore.
5.2 Conclusion
In this paper, we present Audio-Agent, a multimodal framework for both text-to-audio and video-to-audio tasks. Our model offers a conversation-based method for audio generation, editing and composition, facilitating audio generation conditioned on multievent complex descriptions. For the video-to-audio task, we propose an efficient method to achieve visual synchronization. Through extensive experiments, we show that our model can synthesize high-fidelity audio, ensuring semantic alignment with input. Additionally, our work takes an initial, significant step toward multi-event long-condition TTA generation which has not been fully explored.
Appendix A Appendix
A.1 Prompt example for TTA task
We provide our prompt instruction in Table 7 and in context examples in Tables 8 and 9.
@@ A.1 Prompt example for TTA task
A.2 Prompt example for VTA task
@@
A.3 Complex captions for TTA task

### v2

@@ 3.5 Conditional Audio Generation
$\begin{split}\!\!\!\!\!\!\mathbf{Z}^{new}=\text{Softmax}(\frac{\mathbf{Q}\mathbf{K}_{txt}^{\top}}{\sqrt{d}})\mathbf{V}_{txt}+\text{Softmax}(\frac{\mathbf{Q}(\mathbf{K}_{vis})^{\top}}{\sqrt{d}})\mathbf{V}_{vis}\\ \noindent\text{where}\ \mathbf{Q}=\mathbf{Z}\mathbf{W}_{txt}^{q},\mathbf{K}_{txt}=\boldsymbol{c}_{txt}\mathbf{W}_{txt}^{k},\mathbf{V}_{txt}=\boldsymbol{c}_{txt}\mathbf{W}_{txt}^{v},\\ \mathbf{K}_{vis}=\boldsymbol{c}_{vis}\mathbf{W}_{vis}^{k},\mathbf{V}_{vis}=\boldsymbol{c}_{vis}\mathbf{W}_{vis}^{v}\end{split}$ (2)
@@
$L_{\text{simple}}=\mathbb{E}_{\boldsymbol{x}_{0},\boldsymbol{\epsilon},\boldsymbol{c}_{txt},\boldsymbol{c}_{vis},t}\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\big{(}\boldsymbol{x}_{t},\boldsymbol{c}_{txt},\boldsymbol{c}_{vis},t\big{)}\|^{2}.$ (4)
Compared to IP-Adapter [50], our method introduces the video modality into audio generation. Furthermore, since the semantic tokens already incorporate temporal information of the video, we do not need to train an extra timestamp detection module as done by FoleyCrafter [52] to achieve temporal alignment.
@@ 3.6 Implementation Details
For fine-tuning Gemma-2B-it, we set LoRA rank and alpha to be 64 with dropout to be 0.05. We separately train and fine-tune Gemma-2B-it, the projection layers and the cross-attention layers on the AVSync15 [51] datasets. The training and evaluation are conducted on NVIDIA GeForce RTX 4090. Following [50], we set the $\lambda$ to be 0.5 as default.
@@ 4.1 Datasets
Our framework is training-free for the TTA task. The evaluation of our complex generation ability is conducted on AudioCaps [22] dataset. We randomly choose either one caption from the test set or concatenate two of them together with the clause “followed by.” To achieve the optimal solution, the model needs to decompose the input into atomic text instructions, not only limited to two chunks. To better compare with other models, we limit our generation length to the standard 10 seconds. Following [48], we randomly selected 20 captions from each category for the generation. Additionally, to demonstrate Audio-Agent’s ability to make inferences based on complex text conditions, we ask GPT to generate additional long event descriptions containing a series of events without explicitly describing the sound, such as “A man enters his car and drives away”. The number of complex captions is also 20. The baseline methods include AudioGen-v2-medium [24], AudioLDM2-large [29], Auffusion [48], Make-An-Audio 2 [17] and WavJourney [30].
@@
We use AVSync15 for training and evaluation on VTA task. AVSync15 is a curated dataset from VGGSound Sync [3] that has 1500 high video-audio alignment pairs, which is ideal for training and demonstrating temporal alignment between video and audio. Same experiment setting as [52] is used. To better facilitate evaluation, we include some audio generation results in the supplementary material.
@@
AudioGen [24] 49.34% 44.76% 23.98%
AudioLDM2 [29] 47.04% 36.03% 23.33%
Auffusion [48] 50.91% 45.90% 14.40%
Make-An-Audio 2 [17]  55.49% 44.01% 23.73%
WavJourney [30] 34.83% 24.90% 23.71%
Ours (with volume control) 55.26% 48.83% 21.10%
@@
Table 3: Quantitative evaluation on semantic alignment and audio quality. Specifically, Audio-Agent achieves on par performance versus state-of-the-art models in terms of Mean KL Divergence (MKL) [19], CLIP [43] and FID [14] on AVSync15  [51]. See text for explanation on the subpar performance.
@@
SpecVQGAN (Inception) [19] 5.339 6.610 114.44
SpecVQGAN (ResNet) [19] 3.603 6.474 75.56
Diff-Foley [31] 1.963 10.38 65.77
Seeing and Hearing [47] 2.547 2.033 65.82
FoleyCrafter [52] 1.497 11.94 36.80
@@
Table 4: Quantitative evaluation in terms of temporal synchronization. We report onset detection accuracy (Onset ACC) and average precision (Onset AP) for the generated audios on AVSync [51], which provides onset timestamp labels for assessment, following previous studies  [31, 46].
@@
SpecVQGAN(Inception) [19] 16.81 64.64
SpecVQGAN(ResNet) [19] 26.74 63.18
Diff-Foley [31] 21.18 66.55
Seeing and Hearing [47] 20.95 60.33
FoleyCrafter [52] 28.48 68.14
@@
4.2 Evaluation Metrics
The evaluation metrics are summarized as follows: For the VTA task, we use the Frechet audio distance (FAD) to evaluate audio fidelity. Additionally, we utilize the MKL metric [19] and CLIP similarity [43] for audio-video relevance. Furthermore, to evaluate the synchronization of the generated audio in the video-to-audio setting, we use the same evaluation metrics as CondFoleyGen [12], namely # Onset Accuracy, and Onset AP. For the TTA task, we use CLAP similarity [44]
4.3 Evaluation and comparison
Audio-Agent outperforms other baseline methods on the majority of TTA experiment settings, only lagging behind Make-an-Audio 2 by a small margin on the Single Caption setting, which we attribute to the factor of our generation backbone, since we did not train our diffusion model on complex caption datasets. The result can be found in Table 2. Moreover, our method outperforms the original Auffusion backbone model by a significantly increasing margin as the text condition becomes longer and more complex. From the visualization examples in Table 5, we can see our method achieves more fine-grained distinction among events with better control on generation length. Specifically, we notice that with a longer text condition, all of AudioGen [24], AudioLDM2 [29], Auffusion [48] and Make-an-Audio 2 [17] exhibit miss-out events. For example, if the text condition is multi-event such as “Pigeons cooing and bird wings flapping as footsteps shuffle on paper followed by motor sounds with male speaking”, all the above baseline methods fail to generate the motor sound at the end of the audio clip during evaluation. In addition, during our experiments, we found that WavJourney often struggles to generate audio with the precise duration specified by the user, even when the duration is explicitly included in the prompt. Our method avoids these issues by using GPT-4 as a coordinator for effective caption analysis and generation planning, enabling finer distinctions between events. In addition, we found that if the text description contains too many events, the GPT-4 tends to lower the volume of every event to prevent possible volume explosion. Although this improves the sound quality from the human ear, the distinction between each event becomes less clear, which may result in a decrease in CLAP score.
We also notice a significant drop for all these methods on complex captions, since none of them has been trained on this type of text condition which is closer to real-world situations, since it does not require meticulous and explicit descriptions of the characteristics of the sound, but rather describes the scenario for sound generation, thus offering more flexibility for the sound producer. In the supplementary materials, we provide examples of complex results used for evaluation, along with additional comparisons of multi-turn generation and editing.
For the VTA task, our method achieves better visual-audio synchronization compared to other baseline methods, while subpar compared with the current state-of-the-art method in terms of generation audio quality, see Tables 3 and 4. This is indeed expected, as most of the other baseline methods have been trained on multiple larger datasets. We also find that the temporal connector may negatively affect the generated audio quality on a small scale. However, for the evaluation of synchronization, we noticed a significant improvement after the temporal connector was applied, especially for the Onset AP. Without explicit training of a timestamp detector, our method achieves a better performance across all baseline methods in terms of onset Acc and Onset AP, see Figure 6 for illustration.
Figure 5: Comparison with baseline on TTA. From top: our Audio-Agent, Auffusion, Make-An-Audio 2, and WavJourney. To demonstrate audio generation based on long complex text conditions, we ask the respective model to generate audio clips for 20 seconds. The text condition is drawn from the Two Captions category of Table 2: (A) A river stream of water flowing followed by typing on a computer keyboard; (B) A woman delivering a speech followed by a male speech and statics; (C) A vehicle engine revving then accelerating at a high rate as a metal surface is whipped followed by tires skidding followed by a door shutting and a female speaking; (D)  Continuous white noise followed by a vehicle driving as a man and woman are talking and laughing. Observe (by listening to the corresponding audio files in the supplementary material) our method successfully generates multi-event audio at different times based on descriptions, while other baselines mix the generated audio. Particularly for WavJourney, although the boundary between events is more distinct, it consistently does not obey the prompt to generate 20 seconds of audio. The generated audio clips are respectively of length 20 seconds, 30 seconds (truncated here), 16 seconds, and 12 seconds.
Figure 6: Comparison with baseline for VTA generation task. Compared to the baseline, the event occurrence is more explicit. Our method can produce audio that is more aligned and better synchronized with the input video. The corresponding audio files can be found in the supplementary materials.
@@ 4.4 Ablation Studies
We include our ablation study on different LoRA rank values during LLM fine-tuning, see Tables 5 and 6. We find that an increase in trainable parameters sometimes does not necessarily improve the result. Notwithstanding, for a fair comparison, we use the rank value of 60 across all metrics. Additionally during training, we find that the training of the cross-attention layer can converge within 20,000 steps. We notice that the loss curve is not a reliable indicator of the model’s performance. The model can achieve a good performance even when the loss curve remains flat.
Additionally, our framework can be easily integrated with different LLMs for action planning and conversation, apart from GPT-4. In the supplementary material, we included the comparison between various models such as Llama [40], Gemma [39], and Mixtral [20]. With this, our framework’s performance can be improved with the development of LLMs in more precise instruction-following with more sophisticated prompts.
Table 5: Ablation study on AVSync15 dataset with different LoRA rank for semantic alignment and audio quality. During experiments, we keep the value of alpha the same as the rank. See text for further explanation.
@@
Table 6: Ablation study on AVSync15 dataset with different LoRA rank in terms of temporal synchronization. During experiments, we keep the value of alpha the same as the rank. See text for further explanation.
@@ 5 Conclusion and Discussion
Limitation and Future Work Our framework experiences a performance drop when text conditions becomes increasingly complex for the TTA task, where the drop is more severe in other baseline methods. We believe it is a worthwhile direction in the future for understanding long and complex captions with improved fine-grained delineations among multiple and overlapping events. We may also utilize the LLM’s versatility involving audio captioning tasks and video captioning tasks. The above are worthwhile future directions to explore.
Conclusion In this paper, we present Audio-Agent, a multimodal framework for both text-to-audio and video-to-audio tasks. Our model contributes a subbstantial first attempt in offering a conversation-based method for audio generation, editing and composition, facilitating audio generation conditioned on multievent complex descriptions. For the video-to-audio task, we propose an efficient method to achieve visual synchronization. Through extensive experiments, we show that our model can synthesize high-fidelity audio, ensuring high semantic alignment with input. Moreover, our work represents an initial and significant step toward multi-event long-condition TTA generation which has not been fully explored.
Supplementary Material
6 supplementary material
6.1 Prompt example for TTA task
We provide our prompt instruction in Table 7, the volume control variation in Table 10, and in context examples in Tables 8,  911 and  12.
@@ 6.1 Prompt example for TTA task
Table 10: Our prompt instruction for TTA generation
1. A muddled noise of broken channel of the TV.
2. A person is turning a map over and over.
3. Several barnyard animals mooing in a barn.
4. An office chair is squeaking.
5. A flying bee is buzzing loudly around an object.
6. Thunder claps far in the distance.
7. Something goes round that is playing its song.
8. A paper printer is printing off multiple pages.
10.A person crunches through dried leaves on the ground.
**Instructions:**
4. **Requirement**:
4.2. You should include the volume for each generation call in dB following LUFS standard.
4.3. If the user input has multiple events or asks to combine multiple description together, you should have overlapping audios happening in the same range of time. There should have less than three audios in the same time. Overlapping means one audio having smaller start_time than another audios end_time. You should correspondingly adjust the volume for each overlapping calls.
**Response Format:**
- You should only respond in JSON format, following this template:
‘‘‘json
Table 11: Our in-context examples for TTA generation.
**Examples:**
**Example 1:**
- **Your Output**:
‘‘‘json
"plan": "1. Auffusion.generate(’A clap of thunders.’,start_time=2,end_time=5,volume=-10); 2. Auffusion.generate(’Rain pouring outside.’,start_time=0, end_time=10,volume=-30)"
**Example 2:**
- **Your Output**:
‘‘‘json
"plan": "1. Auffusion.generate(’A motor buzzing and humming’,start_time=0,end_time=10,volume=-25); 2. Auffusion.generate(’A man speaking.’,start_time=3,end_time=6,volume=-15)"
**Example 3:**
- **Your Output**:
‘‘‘json
"plan": "1. Auffusion.generate(’A series of machine gunfire.’,start_time=0,end_time=4,volume=-10); 2. Auffusion.generate(’Two gunshots firing.’,start_time=4,end_time=6,volume=-10); 3. Auffusion.generate(’A jet aircraft flies.’,start_time=0,end_time=6,volume=-15); 4. Auffusion.generate(’Soft music playing.’,start_time=6,end_time=10,volume=-20)"
Table 12: Our in-context examples for TTA generation (continue).
**Example 4:**
- **Your Output**:
‘‘‘json
"plan": "1. Auffusion.generate(’Sound of a basketball bouncing on the court.’,start_time=0, end_time=7,volume=-15); 2. Auffusion.generate(’A ball hit the basket’,start_time=5, end_time=7,volume=-10); 3. Auffusion.generate(’People cheering and shouting.’,start_time=7, end_time=10,volume=-10)"
- **Your Output**:
‘‘‘json
"plan": "1. Auffusion.generate(’Sound of a table tennis ball bouncing on the table.’,start_time=0,end_time=7,volume=-10); 2. Auffusion.generate(’People cheering and shouting.’,start_time=7,end_time=10,volume=-10)"
6.2 Prompt example for VTA task
@@
6.3 Complex captions for TTA task
@@
6.4 Comparison with different LLM for TTA task
We present results using different LLMs for generation planning in Table 15. We found that most LLMs can follow the instruction and output atomic descriptions, except Gemma2-27B and Llama-3.1-8B output contains repetitive or unnecessary actions. With stronger backbone, our framework could achieve improved results.
@@ 1. A man enters his car and drives away
LLM Turn 1 Decomposition Results
GPT-4-Turbo 1. ’Car door opening and closing.’, start_time=0, end_time=3; 2. ’Car engine starts.’, start_time=2, end_time=5 3. ’Car driving away.’, start_time=4, end_time=10
Gemma2-27B 1. ’A car door opening and closing.’, start_time=0, end_time=2 2. ’A man closing car door.’, start_time=2, end_time=4; 3. ’A car engine starting.’, start_time=4, end_time=5 4. ’A car driving away.’, start_time=5, end_time=10
Llama-3.1-8B 1. ’A car engine starts and the car drives away.’,start_time=0,end_time=5 2. ’A man enters his car and drives away.’,start_time=0,end_time=5 3. ’A man is talking on the phone.’,start_time=5,end_time=10
Llama-2-7B 1. ’A car door opens.’,start_time=0,end_time=1 2. ’A man gets in the car.’,start_time=1,end_time=3 3. ’Car engine starts.’,start_time=3,end_time=5 4. ’Car drives away.’,start_time=5,end_time=10
Mixtral-8x7B-v0.1 1. ’Car door opening.’,start_time=0, end_time=3 2. ’A man getting into the car.’,start_time=1, end_time=4 3. ’Car door closing.’,start_time=3, end_time=5 4. ’Car engine starting.’,start_time=4, end_time=6 5. ’Car driving away.’,start_time=6, end_time=10
Meta-Llama-3-70B 1. ’A car door opens and closes.’,start_time=0,end_time=2 2. ’A person gets in a car and settles in.’,start_time=0,end_time=2 3. ’A car engine starts.’,start_time=2,end_time=4 4. ’A car drives away.’,start_time=4,end_time=10
Table 15: Comparison of Using Different LLMs on the caption ”A man enters his car and drives away.”.